OpenStudy (anonymous):

3 divided by BLANK equals 12. What is blank?

OpenStudy (anonymous):

I probably don't ask the hardest questions, but I will really appreciate the help! Thanks!

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

3/x=12. 12x=3 3/12= 1/4:)

OpenStudy (anonymous):

Thanks! (: